if a*sin^(-1)x – b*cos^(-1)x=c then
a*sin^(-1)x + b*cos^(-1)x=???? 

Grade:12th pass

3 Answers

Rinkoo Gupta
9 years ago
replace x by -x in given eq.
asin-1(-x) -bcos-1(-x) =c
-asin-1(x)-bcos-1(x)=c
asin-1(x)+bcos-1(x)= -c
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty
Rohit Falake
9 years ago
but options are
  1. 0
  2. [(pi)ab+c(b-a)]/(a+b)
  3. pi/2
  4. [(pi)ab+c(a-b)]/(a+b)
Yash Gupta
5 years ago
asin^-1 x - bcos^-1 x = c 
We have  
bsin^-1 x + bcos^-1 x = b(pi)/2
(a+b)sin^-1 x 
  = b(pi)/2 + c
sin^-1 x =( b(pi)/2 + c)/a+ b
cos^-1 x= [(pi)ab + c(a-b)]/a+b